# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ 2.3 Characterization via differential inclusions
We now present a characterization of solutions to the dual problem (dBPDN) in terms of an initial value problem involving differential inclusions. Let $V\colon\mathbb{R}^{m}\times[0,+\infty)\times\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ denote the objective function in (dBPDN), parameterized in terms of $t$ and $\boldsymbol{b}$:
@@
$\frac{d\boldsymbol{p}}{d\tau}(\tau)\in-\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b}).$ (9)
Differential inclusions generalize the concept of ordinary differential equations to multi-valued maximal monotone mappings [3]. These mappings include, as a special case, the subdifferentials of proper, lower semicontinuous and convex functions (see Definition A.12). Their corresponding differential inclusions are sometimes called gradient inclusions [13, Section 1.3.7, page 10] because they generalize the classical concept of gradient systems [^1] Note that $V(\cdot;t,\boldsymbol{b})$ is differentiable at $\boldsymbol{p}(\tau)\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})\iff\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})=\{\nabla_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})\}$, in which case $\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\nabla_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})$.. Indeed, similarly to how trajectories of gradient systems converge to their critical points (if any exist), the trajectories of gradient inclusions converge to their critical points (if any exist). This fact follows from a variational principle called the minimal selection principle.
@@ 3 The minimal selection principle
Theorem 3.1 (Existence and uniqueness of solutions to differential inclusions with maximal monotone mappings).
Let $g\colon\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semicontinuous, and convex function and let $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial g$. Consider the differential inclusions
$\frac{d\boldsymbol{p}}{d\tau}(\tau)\in-\partial g(\boldsymbol{p}(\tau)),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}.$ (10)
Then, there exists a unique solution $\boldsymbol{p}\colon[0,+\infty)\to\mathrm{dom~{}}\partial g$ satisfying (10). Moreover:
@@
    For every $\tau\geqslant 0$, we have $\frac{d\boldsymbol{p}}{d\tau}(\tau)=\lim_{\begin{subarray}{c}h\to 0\\ h>0\end{subarray}}\frac{\boldsymbol{p}(\tau+h)-\boldsymbol{p}(\tau)}{h}$. Moreover, $\tau\mapsto\frac{d\boldsymbol{p}}{d\tau}(\tau)$ is right-continuous and $\tau\mapsto\left\|{\frac{d\boldsymbol{p}}{d\tau}(\tau)}\right\|_{2}$ is nonincreasing.
@@
    $\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\mathrm{proj}_{\partial g(\boldsymbol{p}(\tau))}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (11)
@@
  • (iii)
    (Asymptotic limit.) If $g$ achieves its minimum at some point, then $\boldsymbol{p}(\cdot)$ converges to such a point:
    $\lim_{\tau\to\infty}g(\boldsymbol{p}(\tau))=\min_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p})\quad\text{and}\quad\lim_{\tau\to\infty}\boldsymbol{p}(\tau)\in\operatorname*{arg\,min}_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p}).$
@@
See [3, Theorem 1, page 147 and Theorem 1, page 159] for the proofs of (i)–(iii). All proofs rely on the fact that the subdifferential $\partial g$ is a maximal monotone mapping (see Definition A.12) and build on most of the material presented in [3, Chapters 1-3]. ∎
@@
In the language of dynamical systems theory, the initial value problem (11) is called a projected dynamical system [74]. Such systems arise naturally in the theory of variational inequalities [2, Chapter 17], as was first noted in [23] and later proved in [24].
3.1 The minimal selection principle for basis pursuit denoising
In the context of this work, the minimal selection principle implies that the gradient inclusions (9) has a unique, global solution satisfying the initial value problem
$\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ (12)
@@ 3.1 The minimal selection principle for basis pursuit denoising
Moreover, the solution converges asymptotically to the unique solution of (dBPDN) when $t>0$ (a solution of (dBP) when $t=0$). We call the projected dynamical system (12) the slow system and its solution the slow solution. In addition, we write
@@
The minimal selection principle suggests we can compute the solution to (dBPDN) (a solution to (dBP) when $t=0$) by calculating the trajectory of the slow system (12) and its asymptotic limit. To do this, we must compute the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})$ for any $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot,t,\boldsymbol{b})$. We turn to this problem next.
@@ 3.2 Computing the minimal selection
Here, we describe how the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})\equiv-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ of the slow system (12) is computed from a cone projection problem, equivalently a nonnegative least-squares (NNLS) problem. To state this succinctly, we will use submatrix notation: Given $\boldsymbol{p}\in\mathrm{dom~{}}V$, we denote by $\bm{A}_{\mathcal{E}(\boldsymbol{p})}$ the $m\times|\mathcal{E}(\boldsymbol{p})|$ the submatrix of $\bm{A}$ with columns indexed by $\mathcal{E}(\boldsymbol{p})$, we write $\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p})}$ to denote its transpose, and we denote by $\bm{D}_{\mathcal{E}(\boldsymbol{p})}$ the $|\mathcal{E}(\boldsymbol{p})|\times|\mathcal{E}(\boldsymbol{p})|$ submatrix of signs $\bm{D}(\boldsymbol{p})$ with rows and columns indexed by $\mathcal{E}(\boldsymbol{p})$. Finally, we denote by $\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p})}$ the subvector of $\boldsymbol{u}\in\mathbb{R}^{n}$ indexed by $\mathcal{E}(\boldsymbol{p})$. With the notation set, we have the following:
@@
For every $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$, the subdifferential $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is non-empty, closed and convex. Hence the projection of $\boldsymbol{0}$ onto $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ exists and is unique (see Definition A.17). The projection is given precisely the NNLS problem (16) or equivalently by its dual problem, the cone projection problem (14).
@@
The NNLS problem (16) may have more than one optimal solution, but the direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is always unique. A detailed review of cone projection algorithms suitable for (14) can be found in [20]. Algorithms for NNLS problems include active set algorithms with “finite-time” convergence, such as the Lawson–Hanson algorithm [41] (see also [12, 19]) and its (highly underappreciated) generalization, Meyer’s algorithm [47].
3.3 The minimal selection yields a descent direction
Lemma 3.1 shows that the minimal selection $-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})$, equivalently the instantaneous direction $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ of the slow system (12), is calculated from a cone projection problem. The next proposition shows that this characterization implies $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ is a descent direction for $V$ and, crucially, obeys an evolution rule.
@@ 3.3 The minimal selection yields a descent direction
    $V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-V(\boldsymbol{p}_{0};t,\boldsymbol{b})=\Delta\left(t\Delta/2-1\right)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}.$ (19)
@@
    $-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b}).$
@@
See Appendix B.1 for the proof. ∎
Remark 3.3 (Drift to infinity).
Without the full row rank assumption $\mathrm{rank}({\bm{A}})=m$, there may be some $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ for which $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\neq\boldsymbol{0}$ with $\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$. In this case, $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ for every $\Delta\in[0,+\infty)$, hence $\lim_{\Delta\to+\infty}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))=+\infty$ by (18). This will have implications for the calculation of the trajectory of the slow system (12) when $\mathrm{rank}({\bm{A}})\neq m$.
4 Dynamics of the slow system
4.1 Explicit local solution
The evolution rule (20) describes how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ evolves locally in $\Delta>0$ along the line $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. This evolution is local because it holds on some possibly finite interval $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, with $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ given by (18). This behavior, a consequence of the minimal selection principle, fundamentally characterizes the dynamics of the slow system (12). It also appears as the main obstacle for computing the slow solution.
Here, we show that this behavior is no obstacle as it appears: Locally, the slow system evolves as that of its non-projected counterpart, in the sense that
$\frac{d\boldsymbol{p}(\tau)}{d\tau}=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0})=-t(\boldsymbol{p}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\ \text{ for small enough times $\tau>0$.}$
Its local solution can therefore be computed explicitly, as the following Theorem makes precise.
@@ 4.1 Explicit local solution
$\frac{d\boldsymbol{p}}{d\tau}(\tau)=e^{-t\tau}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\quad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (21)
@@
In particular, the slow solution is given explicitly on $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by
@@
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)=-t(\boldsymbol{p}_{a}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\qquad\boldsymbol{p}_{a}(0)=\boldsymbol{p}_{0}.$ (25)
satisfies the slow system (12) on $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ and conclude using uniqueness.
First, consider the affine system (25). A short calculation shows that its unique, global solution is given by
@@
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)$ $=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (26)
@@
$-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (27)
whenever $0\leqslant f(\tau,t)<\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Notice $\tau\mapsto f(\tau,t)$ increases monotonically from 00 to $1/t$ (0 to +$\infty$+ ∞) when $t>0$ ($t=0$). Hence if $t=0$, the largest value of $\tau$ for which (27) holds is $\Delta_{*}(\boldsymbol{p}_{0};0,\boldsymbol{b})$. If $t>0$ and $01-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>$ 0, then $(1-e^{-t\tau})/t$ is equal to $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ at $-\ln\left(1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})\right)/t$. If $t>0$ and $1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\leqslant 0$, then (27) holds for every $\tau\geqslant 0$. Taken together, we find
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})\ \text{for every $\tau\in[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$}.$

### v2

@@ 2.3 Characterization via differential inclusions
We now present a characterization of solutions to the dual problem (dLASSO) in terms of an initial value problem involving differential inclusions. Let $V\colon\mathbb{R}^{m}\times[0,+\infty)\times\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ denote the objective function in (dLASSO), parameterized in terms of $t$ and $\boldsymbol{b}$:
@@
$\dot{\boldsymbol{p}}(\tau)\in-\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b}).$ (9)
Differential inclusions generalize the concept of ordinary differential equations to multi-valued maximal monotone mappings [3, 8]. As a special case, these mappings include subdifferentials of proper, lower semicontinuous and convex functions [43, Theorem 12.17]. Their corresponding differential inclusions are called gradient inclusions [10] because they generalize the classical concept of gradient systems. Indeed, similarly to how trajectories of gradient systems converge to their critical points (if any exist), trajectories of gradient inclusions converge to their critical points (if any exist). This fact follows from a variational principle called the minimal selection principle.
@@ 3 The minimal selection principle
Theorem 3.1 (Existence and uniqueness of solutions to gradient inclusions).
Let $g\colon\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semicontinuous, and convex function and let $\boldsymbol{p}_{0}\in\mathrm{dom~}\partial g$. Consider the gradient inclusions
$\dot{\boldsymbol{p}}(\tau)\in-\partial g(\boldsymbol{p}(\tau)),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}.$ (10)
Then, there exists a unique solution $\boldsymbol{p}\colon[0,+\infty)\mapsto\mathrm{dom~}\partial g$ satisfying (10). Moreover:
@@
    The function $\tau\mapsto\dot{\boldsymbol{p}}(\tau)$ is right-continuous and the function $\tau\mapsto\left\|{\dot{\boldsymbol{p}}(\tau)}\right\|_{2}$ is nonincreasing.
@@
    If $g$ achieves its minimum at some point, then $\boldsymbol{p}(\cdot)$ converges to such a point:
    $\lim_{\tau\to+\infty}g(\boldsymbol{p}(\tau))=\min_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p})\quad\text{and}\quad\lim_{\tau\to+\infty}\boldsymbol{p}(\tau)\in\operatorname*{arg\,min}_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p}).$
  • (iii)
@@
    $\dot{\boldsymbol{p}}(\tau)=-\mathrm{proj}_{\partial g(\boldsymbol{p}(\tau))}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (11)
@@
See [3, Theorem 1, page 147 and Theorem 1, page 159] for the proofs of (i)–(iii). ∎
@@
In the language of dynamical systems theory, the initial value problem (11) is called a projected dynamical system [19]. Such systems arise naturally in the theory of variational inequalities [2, Chapter 17], as was first noted in [18].
3.1 The minimal selection principle for the lasso problem
In the context of this work, the minimal selection principle implies the system of gradient inclusions (9) has a unique, global solution satisfying the initial value problem
$\dot{\boldsymbol{p}}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ (12)
@@ 3.1 The minimal selection principle for the lasso problem
Moreover, the solution converges asymptotically to the unique solution of (dLASSO) when $t>0$ (a solution of (dBP) when $t=0$). The projected dynamical system (12) is called the slow system and its solution the slow solution. In addition, we will write
@@
The minimal selection principle suggests we can compute a solution to (dLASSO) or (dBP) by calculating the asymptotic limit of the slow system (12). To do this, we must compute the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})$ for any $\boldsymbol{p}_{0}\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot,t,\boldsymbol{b})$. We turn to this problem next.
@@ 3.2 Computing the minimal selection
Here, we describe how the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})\equiv-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ of the slow system (12) is computed from a cone projection problem or, equivalently, a nonnegative least-squares (NNLS) problem. To state this succinctly, we will use submatrix notation: Given $\boldsymbol{p}\in\mathrm{dom~}V$, we denote by $\bm{A}_{\mathcal{E}(\boldsymbol{p})}$ the $m\times|\mathcal{E}(\boldsymbol{p})|$ the submatrix of $\bm{A}$ with columns indexed by $\mathcal{E}(\boldsymbol{p})$, we write $\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p})}$ to denote its transpose, and we denote by $\bm{D}_{\mathcal{E}(\boldsymbol{p})}$ the $|\mathcal{E}(\boldsymbol{p})|\times|\mathcal{E}(\boldsymbol{p})|$ submatrix of signs $\bm{D}(\boldsymbol{p})$ with rows and columns indexed by $\mathcal{E}(\boldsymbol{p})$. Finally, we denote by $\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p})}$ the subvector of $\boldsymbol{u}\in\mathbb{R}^{n}$ indexed by $\mathcal{E}(\boldsymbol{p})$. With the notation set, we have the following:
@@
For every $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$, the subdifferential $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is non-empty, closed and convex. Hence the projection of $\boldsymbol{0}$ onto $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ exists and is unique (see Definition A.17). The projection is given precisely the NNLS problem (16) or equivalently by its dual problem, the cone projection problem (14). Equation (17a) states the classical Karush-Kuhn-Tucker (KKT) conditions applied to the NNLS problem (16). To obtain equation (17b), we write
@@
Problem (16) may have more than one optimal solution, but the direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is always unique. A review of cone projection algorithms can be found in [15]. Algorithms for NNLS problems include active set algorithms with “finite-time” convergence, such as the Lawson–Hanson algorithm [32] and its generalization, Meyer’s algorithm [36].
4 Dynamics of the slow system
4.1 The minimal selection is a descent direction
Lemma 3.1 shows that the instantaneous direction $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ of the slow system (12) can be calculated from a cone projection problem. The following proposition shows that this characterization implies $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ is a descent direction for $V$ and, crucially, obeys an evolution rule.
@@ 4.1 The minimal selection is a descent direction
    $V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-V(\boldsymbol{p}_{0};t,\boldsymbol{b})=-\Delta\left(1-t\Delta/2\right)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}.$ (19)
@@
    $(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b}).$
@@
See Appendix B.1. ∎
4.2 Explicit local solution of the slow system
The evolution rule (20) describes how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ evolves locally in $\Delta>0$ along the line $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. This evolution is local because it holds on some possibly finite interval $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, with $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ given by (18). Here, we use this evolution rule to show that the slow system evolves as that of its non-projected counterpart, in the sense that
$\dot{\boldsymbol{p}}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0})\equiv-t(\boldsymbol{p}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\ \text{ for small enough times $\tau>0$.}$
Its local solution can therefore be computed explicitly, as the following Theorem makes precise:
@@ 4.2 Explicit local solution of the slow system
$\dot{\boldsymbol{p}}(\tau)=e^{-t\tau}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\quad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (21)
@@
In particular, the slow solution is given explicitly on $0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by
@@
$\dot{\boldsymbol{p}}_{a}(\tau)=-t(\boldsymbol{p}_{a}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\qquad\boldsymbol{p}_{a}(0)=\boldsymbol{p}_{0}.$ (25)
satisfies the slow system (12) on $0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ and conclude using uniqueness. First, consider the affine system (25). A short calculation shows that its unique, global solution is given by
@@
$\dot{\boldsymbol{p}}_{a}(\tau)$ $=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (26)
@@
$\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (27)
whenever $0\leqslant f(\tau,t)<\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Notice $\tau\mapsto f(\tau,t)$ increases monotonically from 0 to $1/t$ (0 to $\infty$) when $t>0$ ($t=0$). Hence if $t=0$, the largest value of $\tau$ for which (27) holds is $\Delta_{*}(\boldsymbol{p}_{0};0,\boldsymbol{b})$. If $t>0$ and $1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0$, then $(1-e^{-t\tau})/t$ is equal to $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ at $\tau=-\ln\left(1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})\right)/t$. If $t>0$ and $1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\leqslant 0$, then (27) holds for every $\tau\geqslant 0$. Taken together, we find
$\dot{\boldsymbol{p}}_{a}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})\ \text{for every $\tau\in[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$}.$